OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions?

OpenStudy (cherylim23):

For solutions to be rational, discriminant must be zero, or a perfect square. For solutions to be rational AND identical, discriminant must be zero.

OpenStudy (anonymous):

so the answer is when the radicand equals zero

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

if the radicand (aka discriminant ) \(b^2-4ac=0\) then the quadratic formula tells you the zeros are \(-\frac{b}{2a}\) aka the vertex. it also means you have a perfect square, like say \((x-4)^2=0\)